OpenStudy (anonymous):

Prove there are no positive integer solutions to the equation a^2 - b^2 = 10

OpenStudy (amistre64):

a^2 = b^2 - 10 would be my first instinct

OpenStudy (ash2326):

a^2=b^2+10

OpenStudy (amistre64):

yeah, that one ;)

OpenStudy (amistre64):

1,4,9,16,25,36 are your squares

OpenStudy (ash2326):

a and b either be positive or negative doesn't matte

OpenStudy (amistre64):

11, 14, 19, 26, 46, 59, ... would be the resulting sequence if we can work with it

OpenStudy (ash2326):

Amistre64 but there is no end to the square list it can be 64,81and so on also. we gotta prove that a be an irrational

OpenStudy (amistre64):

right, was thinking maybe a proof by induction mighta been in there someplace ... how bout, since a and b are to be positive integers, then a,b not= 0 a^2=b^2+10 1=b^2/a^2 + 10/a^2 (b/a)^2 = -10/a^2 b/a = sqrt(-10/a^2) b/a = sqrt(-10)/a

OpenStudy (ash2326):

u skipped a 1 in the third step

OpenStudy (amistre64):

its monday right? ;)

OpenStudy (anonymous):

\[(a+b)(a-b)=10\] \[a+b=5\] \[a-b=2\] \[a=\frac{7}{2}\] nope \[a+b=10\] \[a-b=1\] \[a=\frac{11}{2}\] nope

OpenStudy (anonymous):

if a in an integer and b is an integer then \[a+b\], \[a-b\] are both integers. and there are only two ways to factor 10 as the product of two integers, namely \[10=5\times 2\] or \[10=10\times 1\] since neither way gives integer a and b that is it

OpenStudy (anonymous):

i guess i should add i am using the fact that \[a^2-b^2=10\iff (a+b)(a-b)=10\]

OpenStudy (amistre64):

\[\frac{a^2}{10}-\frac{b^2}{10}=1\] is a hyperbola that fer one

OpenStudy (amistre64):

i would say that from this: \[ a \ne b\]\[a>b\]narrows down any proving

OpenStudy (amistre64):

and of course satellites integer factors of 10 .... i feel a bit dense today lol